OpenStudy (anonymous):

Determine the angle x in the triangle given below with AC = 2 and BC = 13. ( it is a right angles triangle: CB is the hypotenuse; AC is where C^AB is where the 90 degreee angle is and C^BA is x)

OpenStudy (anonymous):

easier to label triangles as angles A, B, C and opposite sides a, b, c respectively.

OpenStudy (anonymous):

sin(x)=2/13 and so find the inverse of sin over 2/13 for the angle x.

OpenStudy (anonymous):

find BA first whichbappears to be the height if am not wrong h=sqrt of 169(13^2)-4(2^2) h=sqrt of 165 h is approximately 12.8452 cos theatre = adj/hyp =cos inverse of 12.8452/13 cos-1 of 0.14563 =81.6 degrees

OpenStudy (anonymous):

or post the picture because you would have an answer in a nanosecond

OpenStudy (anonymous):

im trying to :-)

OpenStudy (anonymous):

mathfrick: i think x can't be 81.6 degrees, because the side opssite to x is just 2 units in lenght.

OpenStudy (anonymous):

can that be opened?

OpenStudy (anonymous):

can anyone see the pic I attached?

OpenStudy (anonymous):